# Unveiling Topological Structures in Text: A Comprehensive Survey of Topological Data Analysis Applications in NLP

### v1

@@ 5.4 Improvement in TDA Feature Extraction
Unlike, some other data modalities that have a distinct shape, texts take the shape of their numerical representation. This means that the shape of texts represented using TF-IDF will embody the TF-IDF representation, which tends to be different from other embeddings, such as the Transformer Pre-trained embeddings. While, the difference can be supported by the different linguistic features they capture such as semantic, syntactic, and lexical features, the differences in shape can be un-intuitive, making it difficult to determine the best numerical representation to extract the most useful TDA features. Therefore, we must find novel ways of representing texts numerically, for example, the use of symbolic representations or better ways to use the numerical representations that exist in such a way that is advantageous for extracting the best TDA features.
@@ 5.5 Adversarial Robustness of TDA Features
Robustness to noise, particularly adversarial perturbations, has been an important research topic in the NLP community. While it is encouraging that such robustness of TDA features is promising, there have been only a few works in this direction Perez and Reinauer (2022); Chauhan and Kaul (2022). For instance,  Perez and Reinauer (2022) show that their topologically-augmented BERT model is much more robust than the base BERT model when tested against perturbations generated by TextAttack Morris et al. (2020)Chauhan and Kaul (2022) also shows that there are some weak correlations between persistent homology features of a trained BERT model and its adversarial robustness against several state-of-the-art attack methods. Nevertheless, some results confirm that TDA is more robust, while others were inconclusive. Additionally, all existing works only evaluate on BERT model with simple attack mechanism, missing other security scenarios such as poisoned or backdoor attacks. Therefore, we call for a more comprehensive robustness evaluation of TDA on NLP models, especially from both NLP, ML and security community.
@@ 5.6 Topological Deep Learning for NLP
Due to the benefits of TDA and deep learning, a new niche field is born - Topological Deep Learning (TDL). Initially, TDL is described as an ensemble of topological features extracted by TDA techniques such as persistent homology and deep learning features. However, as the field has advanced, a new definition for TDL has emerged - “the collection of ideas and methods related to the use of topological concepts in deep learning” Papamarkou et al. (2024). This allows a deep learning model to be integrated more deeply with concepts of algebraic topology, such as the introduction of simplicial neural networks (NN), which are NNs with layers made up of simplicial complexes. This deeper integration of TDA into NNs makes TDL particularly useful for the explosion of data we are currently experiencing, especially because these datasets are now more often high-dimensional.
These high-dimensional data need better tools for processing as the current tools shrink the dimension, encouraging loss of information. In NLP, one particular approach to integrate TDA with high-dimensional NLP embeddings has been the utilization of text in graphical forms, which have been shown to yield better results than directly using texts as a sequence of tokens Zhong et al. (2020); Liu et al. (2023); Phan et al. (2023). More research is still needed to validate such an approach to a variety of NLP tasks and problems.
@@ 6 Conclusion
We surveyed 85 papers that applied TDA, using theoretical and non-theoretical approaches in NLP. In the theoretical approaches, researchers are more concerned with explaining linguistic phenomena from a topological perspective. For the non-theoretical approaches, several categories can be used for the taxonomy. Some are binary categories, yielding uninformative distinctions on the non-theoretical applications of TDA in NLP, these include - TDA techniques which are Persistent Homology and Mapper; data modality which is text and speech; and learning types which are supervised and unsupervised. While the more informative categories are tasks and numerical representation. Tasks are divided into 7 categories. Next, the main taxonomy for the non-theoretical approaches is the numerical representation used on texts before extracting TDA features. This is because, in NLP the bottleneck for extracting useful TDA features is the quality of the numerical representation techniques. Finally, based on the utility of TDA in NLP, we present the first survey which discusses the novel applications of TDA in NLP, as well as the challenges and open problems observed.
Acknowledgments
Adaku would like to thank Dr. Charlie Dagli for his encouragement, reading the paper drafts, and providing invaluable recommendations.
@@ Appendix A Appendix
Tfemčinas (2018) word sense induction and disambiguation Pre-trained (Word2Vec, GloVe) Unsupervised Text Persistent Homology
@@
Zadrozny (2021a) investigates the manifestations of intelligence and understanding in neural networks Pre-trained (Transformers - GloVe) Supervised Text Persistent Homology
@@
TOPFORMER Uchendu et al. (2024) deepfake text detection Pre-trained (Transformers - pooled_output) Supervised Text Persistent Homology

### v2

@@ 5.5 Improvement in TDA Feature Extraction
Unlike some other data modalities that have a distinct shape, texts take the shape of their numerical representation. However, different numerical representations capture different linguistic features, making it challenging to intuitively interpret their topological differences. Therefore, we must find novel ways of representing texts numerically, for example, the use of symbolic representations or better ways to use the numerical representations that exist in such a way that is advantageous for extracting the best TDA features.
@@ 5.6 Adversarial Robustness of TDA Features
Robustness to noise, particularly adversarial perturbations, has been an important research topic in NLP. While such robustness of TDA features is promising, there have been only a few works in this direction Perez and Reinauer (2022); Chauhan and Kaul (2022). For instance,  Perez and Reinauer (2022) show that their topologically-augmented BERT model is much more robust than the base BERT model when tested against perturbations generated by TextAttack Morris et al. (2020).  Chauhan and Kaul (2022) also shows that there are some weak correlations between persistent homology features of a trained BERT model and its adversarial robustness against several state-of-the-art attackers. However, all existing works only evaluate on BERT model with simple attack mechanism, missing other security scenarios such as poisoned or backdoor attacks. Therefore, we call for a more comprehensive robustness evaluation of TDA on NLP models, especially from the NLP, ML, and security communities.
@@ 5.7 Topological Deep Learning for NLP
Due to the benefits of TDA and deep learning, a new niche field is born - Topological Deep Learning (TDL) a “the collection of ideas and methods related to the use of topological concepts in deep learning” Papamarkou et al. (2024). TDL allows a deep learning model to be integrated more deeply with concepts of algebraic topology, such as the introduction of simplicial neural networks (NNs), which are NNs with layers made up of simplicial complexes. This deeper integration of TDA into NNs makes TDL particularly useful for the explosion of high-dimensional data. These high-dimensional data need better tools for processing as the current tools shrink the dimension, resulting in information loss. In NLP, one particular approach to integrate TDA with high-dimensional NLP embeddings has been the utilization of text in graphical forms, which have been shown to yield better results than directly using texts as a sequence of tokens Zhong et al. (2020); Liu et al. (2023); Phan et al. (2023). Nevertheless, more research is still needed to validate such an approach.
@@ 6 Conclusion
Our world is currently experiencing an explosion of data and an equally explosion of computational techniques to process such data. Machine Learning (ML) is the most popular of these computational methods, however, while its benefits are numerous, it has a few limitations. The biggest of the limitations of ML is its inability to sufficiently process data that is high-dimensional, imbalanced, noisy, and scarce. Therefore, a small community of NLP researchers emerged to tackle this limitation by proposing using TDA to tackle difficult NLP tasks. These researchers employ two TDA techniques - Persistent Homology and Mapper to solve NLP tasks using theoretical and non-theoretical approaches. This yielded 87 papers which we comprehensively surveyed in this paper. Finally, we conclude that while the applications of TDA in NLP have come a long way from 2013 to now, there is still room for improvement, specifically in reducing the barrier of entry for non-TDA experts to apply it to their NLP tasks.
@@ Appendix A Appendix
Temčinas (2018) word sense induction and disambiguation Pre-trained (Word2Vec, GloVe) Unsupervised Text Persistent Homology
@@
Wright and Zheng (2020) distinguish subsets in data Pre-trained (Word2Vec) Unsupervised Text Persistent Homology
@@
Zadrozny (2021a) investigates the manifestations of intelligence and understanding in neural networks Pre-trained (GloVe) Supervised Text Persistent Homology
@@
TOPFORMER Uchendu et al. (2024) deepfake text detection Pre-trained (Transformers - Hidden) Supervised Text Persistent Homology
@@
Sun and Nelson (2023) Correlation between sentence vectors Pre-trained (Word2Vec, Transformers - Hidden) Unsupervised Text Persistent Homology